Break-even analysis:

Suppose the Starbucks pays 30000 TL/month rent. Prepare a breakeven analysis for the Starbucks Cafe. Please justfy your assumptions and your choice of data. Note: total daily expense of this Starbucks cafe 15.000TL (including Utilities and salaries), 6 full time employees, 1700 customers per day.

Alao, daily revenue is 20.000
Answer #1

Below is the breakeven unit formula

Break even quantity = Fixed Cost / (sale value per unit-variable cost per unit)

In this case, Fixed Cost is rent which is TL. 30000 per month

Variable cost to serve one customer is = per day daily expenses/total number of customer visited = 15000/1700 = 8.82 TL per customer

Revenue per customer = total revenue per day / total number of customer in da day= 20000/1700= 11.76

Break even quantity (number of customers) = 30000 /(11.76-8.82) = 10200

Number of days it will take= 10200//1700= 6 days to get the break even for the month.

Assumption: every day customer visit is equal to 1700 and the total revenue is TL 20000.